### v1

# Goal Inference from Open-Ended Dialog

Abstract
We present an online method for embodied agents to learn and accomplish diverse user goals. While offline methods like RLHF can represent various goals but require large datasets, our approach achieves similar flexibility with online efficiency. We extract natural language goal representations from conversations with Large Language Models (LLMs). We prompt an LLM to role play as a human with different goals and use the corresponding likelihoods to run Bayesian inference over potential goals. As a result, our method can represent uncertainty over complex goals based on unrestricted dialog. We evaluate our method in grocery shopping and home robot assistance domains using a text-based interface and AI2Thor simulation respectively. Results show our method outperforms ablation baselines that lack either explicit goal representation or probabilistic inference.
I Introduction
AI agents and robots must quickly learn and carry out many different user tasks in real-time. For example, a home robot assistant may need to adapt to various household preferences and routines. Imagine a scenario where your robot assistant is tasked with gathering ingredients to bake a cake for you. Depending on who you are, you may want different ingredients. If you only want a basic cake, your ideal recipe is eggs, milk, sugar, and flour. However, if you are allergic to gluten, you would want gluten-free flour. If you prefer strawberry cake, then you would want to get strawberries. Every human has a different set of preferences. This level of variation is challenging, if not impossible, for system designers to anticipate in advance. In order for agents/robots to perform or assist with tasks for humans, they must first be able to learn the preferences of the humans that they are trying to help.
To address this challenge, we propose a new method GOOD (Goals fOr Open-ended Dialogue) that combines the best parts of offline and online approaches. It uses Large Language Models (LLMs) to infer natural language representations of user goals. This allows our method to represent a flexible, open-ended set of possible goals during an online interaction. As a result, the method combines the flexibility and representation power of offline preference tuning methods [1, 2] with the data efficiency and uncertainty quantification of online methods that learn rewards based on a set of engineered features [3, 4]. This allows our method to represent uncertainty over goals that may not have been explicitly engineered or anticipated in advance. In order to learn goals efficiently, we use natural language dialog with the user instead of, e.g., best-of-k comparisons present in [5, 6, 7, 8].
However, applying traditional Bayesian inference to natural language goals presents two significant challenges. First, it is intractable to enumerate the space of all possible natural language expressions. To address this, we run Bayesian inference over a reduced set of explicit hypothesized goals. We use LLM modules to maintain this set of goals. Second, running Bayesian inference requires a likelihood function to represent the conditional distribution of a dialog for different candidate goals. Our key insight is that we can leverage an LLM’s ability to role-play as a human with an explicit goal to define this likelihood.
In this paper, we make the following contributions:
  1. 1.
    We propose a Bayesian inference method that can track a distribution over natural language goals given unrestricted dialog with a user;
  2. 2.
    We design a goal management system that tracks an explicit set of plausible goals that can be used for inference; and
  3. 3.
    We demonstrate that this approach can track a wide range of user goals in a grocery shopping assistant and a home robot assistant domain. Our results indicate that Bayesian inference over an explicit representation of goals is a promising approach to flexible alignment of generative agents.
Figure 1: Offline/RLHF methods are data heavy, but is flexible to accommodate many tasks and domains. Online methods are data efficient, however are very domain specific. To accommodate human preferences from conversations, our method uses the best of both worlds and is data efficient and is generalizable for a broad set of tasks and domains.
II Background and Related Work
II-A Preference Learning and NL Probabilistic Reasoning with LLMs
Offline preference tuning methods [1, 2] are data heavy but are generalizable to many domains and tasks. Online methods are data efficient but often task specific [3]. Our method combines both aspects of being data efficient, can be generalized across many domains and tasks, and is online. Previous work [9, 10], show improvement with using LLMs for learning human preferences and NL probabilistic reasoning. Previous methods such as [11, 8] often rely on asking the most informative questions. Our method does not make these unrealistic assumptions on interactions and is flexible to other interaction methods. Unlike existing preference learning techniques that expect structured forms of data like yes/no queries or comparisons ([5, 6, 7, 8]), our method allows for representing goals that the designer did not explicitly engineer in advance.
Previous works such as [12, 13, 14] show that LLM agents can have different roles to achieve tasks. Our work leverages this concept and has different LLM calls that focus on specific tasks in our pipeline to be more efficient with how much information each LLM call has in memory.
II-B Human Preferences or Human Interaction with LLMs in Robotics
Many works incorporate human preferences in doing robotic tasks or generating robotic plans. However, most of these either rely on best-of-k comparison ([7]), or learns rewards based on a specific set of engineered features, making it difficult to generalize to other assistance tasks or scenarios ([15]). Some works involve human-robot interactions that are much more limited or not through active conversations ([16, 17, 4]). Our work leverages natural language goals and LLMs for flexibility and understanding human preferences through language interactions.
III Method
Figure 2: Method Overview Diagram.Method Overview Diagram. The pipeline consists of four main modules: Conversation, Inference, Goal Management, and Action. The process iterates until the task is completed or the conversation limit is reached.
The key idea behind our method is tracking possible human goals and how likely they are as the dialog between the human and the agent goes on. To track the possible space of human goals, we instantiate a finite goal set to which we can add goals if our inference finds that the human preferences in the conversation so far are not represented in the goals, or remove unlikely goals. Given the updated goal set, we infer the likeliest goals and select actions based on them. We continue the conversation rounds until the task is completed.
III-A Preliminaries
Typical Bayesian preference learning methods interpret human input $u$ as evidence for the person’s goal. Given a new input $u$, these methods model the likelihood $P(u\mid g)$ using, for example, models from econometrics and cognitive science, then perform goal belief updates as follows:
$P(g\mid u)=\frac{P(u\mid g)P(g)}{\sum_{\hat{g}\in G}P(u\mid\hat{g})P(\hat{g})}$ (1)
This formulation presents two challenges for open-ended goal inference. First, how should we flexibly represent the goals themselves? Typical methods ([18], [19]) define goals as $x,y$ locations in navigation or continuous parameters trading off features in a reward function, but these approaches restrict the set of goals or preferences that can be learned.
Algorithm 1 Our Method, where $G$ is the goal list, $q$ is the robot query, $u$ is human utterance, $hp$ is the human profile, $s$ is the task status, $t$ is the transcript so far.
1:$G$ = [‘Unspecified‘]
2:while task not complete & convo num $<20$: ## Generate a round of convo $q,u,t$ = Conversation($rt,s,t,hp$) likelihoods = Belief Update(llm, $G,q,u,t$) if argmax[likelihoods] == ’Unspecified’: add($G$) remove($G$) ## update the likelihoods for new set of goals likelihoods = Belief Update(llm, $G,q,u,t$) likely goals = sort(likelihoods)[0:$k$] ## Take actions based on $k$ likely goals action history, complete = Action(llm, likely goals) if complete == True then end else: convo num +=1+=1+ = 1
Next, what should the likelihood $P(u\mid g)$ even be? Prior work has modeled this as the Boltzmann noisily rational model that assumes humans select inputs in proportion to their exponentiated reward ([20], [21]). However, this model is an oversimplification compared to reality because humans can be biased, myopic, or not even be aware of their internal reward ([22], [23]). Finally, how can we keep track of the set of goals $G$ in the denominator? Prior work has system designers define a set of possible goals ahead of time, but thinking of all possible goals in any environment or for any human preference a priori is unreasonable.
Our method consists of four modules. The Conversation Module produces a round of robot query and a human utterance. The Inference Module does inference on the goal list to find the most likely goals. The Goal Management Module is responsible for managing the goal list. The Action Module generates and takes a sequence of actions based on the most likely goals, and checks to see if the task is completed. The overview of the pipeline is shown in 2 and in Pseudocode 1. Our method elegantly tackles all these questions by leveraging natural language and powerful LLM priors.
III-B Conversation Module
We use a LLM to talk to the user about the task. The conversation module enables dialogue between a robot that asks questions (a “robot query”) and for a human that answers those questions (a “human utterance”). To generate a robot query, an LLM is prompted to generate a question given a description of the robot/agent task, transcript of the chat so far, and current status of a task. Our experiments use LLMs to roleplay as a human with a certain “human profile”. To generate a human utterance, an LLM is prompted to generate a response given the robot’s task, a human profile, the current task status, and phrase for the human to respond with if the task is completed. See Section VII for prompt details. The framework is flexible and can be easily adjusted to allow for user text input or other text generation methods.
III-C Inference Module for Belief Space Update
To decide if the robot’s goal list needs to be altered or which action(s) to take, we next implement an Inference Module. The Inference Module assigns probability to each of the tracked goals.
In order to track a Bayesian posterior over a set of goals, we need to calculate the probability of a goal given a human utterance, $P(g|u)$. Using LLMs to calculate $P(g|u)$ according to Equation 1, we need a model for $P(u|g)$ – a model that tells us what utterances a person would choose conditioned on having a particular goal. LLMs are good at roleplaying ([24]) and have been demonstrated to possess strong common sense priors ([25], [26]). We leverage these strong priors to model $P(u\mid g)$ as an LLM query $\pi_{\text{LLM}}(u,g)$. By prompting the LLM to generate utterances given a goal, we can use the model’s output probabilities as an approximation of this likelihood.
$P(u|g)=\frac{\pi_{\text{LLM}}(u,g)}{\sum_{\hat{g}\in G}\pi_{\text{LLM}}(u,\hat{g})}$ (2)
Consider maintaining a belief over two possible goals: “I want cocoa” and ‘I want noodles” given the utterance “I’d like something sweet.” Our method queries an LLM for the likelihood of the utterance based on the prompt ‘Role-play as a person who wants $g$’, where $g$ is replaced with each possible goal.” Given a human utterance is “I want a cake”, the likelihood of “I want cocoa” would be higher than “I want noodles”. The sum of the logits for “I want a cocoa” when the LLM is provided that “I want a cake” is the true goal is higher than the sum of the logits for “I want a cake” when the LLM is provided that “I want noodles” is the true goal. See Section VII for LLM query details.
III-D Goal Management Module
In exact Bayesian inference, we would apply the inference module with all possible goals. This is clearly intractable. As a result, we implement a module for maintaining a set of potential goals to iterate over. We instantiate this list with natural language goals.
To implement this module, we need to be able to propose new goals and remove unlikely goals. To propose new goals, a LLM is prompted to return a list of goals that can be added given the current goal list, the conversation transcript, and the robot’s task. We remove unlikely and unsafe goals through two ways. 1) A LLM is prompted to return a list of goals to remove that should not be taken given the robot’s task and the conversation transcript. 2) If a goal is the least likely from our inference methods for more than $n$ rounds of conversation, then it should be removed from the goal list.
A key challenge for this module is determining when to propose new goals or remove unlikely ones. This is important because LLMs are expensive and slow, so we need to limit number of calls. We approach this problem by comparing the utterance likelihood to the corresponding likelihood if human has no explicit goal. Consider our example from figure 2. Imagine that the human instead says “I’d like something refreshing”. In this case, the agent should trigger the goal editing module to propose a new goal. We do this with an extra hypothesis that simply prompts the language model to role-play as a human, but does not provide an explicit goal. We call this the “Unspecified Goal”. If the “Unspecified Goal” is the most likely hypothesis, we trigger the goal editing module. In our example, ’I’d like something refreshing’ is more likely for the “Unspecified Goal” than either of the alternatives, so the goal editing module is triggered to propose and add goals. The “Unspecified Goal” can be included like any other goal in the goal list during the Inference Module for every round of conversation. The Inference Module is called twice during the pipeline, the first time to decide if goals should be added (if “Unspecified Goal” is the most likely), and the second time to determine the most likely goals for the Action Module.
The goal removal process is called every round of conversation, to remove any irrelevant goals as soon as possible. This helps with reducing the amount of calls to an LLM for unlikely goals as we iterate over the goal list during the Inference Module.
III-E Action Module
The action module is responsible for two things: 1) selecting action(s) that are available and taking them given the likely goals from the inference module, and 2) checking if the task is completed. This module is domain dependent, and can be easily adapted to other domains and planners. Our implementation will be further explained in more detail in the Experiments section.
III-E1 Do Action
The primary purpose of this module is for the robot/agent to take actions. We do this by prompting a language model to pick actions given the most likely goals and information about their likelihood magnitudes. If the most likely goal is “Unspecified Goal”, then none of the non-unspecified goals present accurately represent the human preference, so proceed to do “no action”. If another goal is the most likely, then the current goal list represents the human preference well, so pick the top $k$ goals from the goal list. This is a benefit of our method because the action module can take actions that are useful for multiple potential likely goals. The module maintains a record of the actions that have been taken to provide context for other modules.
The LLM prompt for Do Action includes information about the previous transcript, the list of possible actions, the most likely goal, and the next $k$ likely goals in order. The prompt will return the list of actions that should be taken in order to accomplish the goal. How the list possible actions are constructed is explained more in the Experiments section.
III-E2 Task Completion Check
After a sequence of actions is taken, we need to check if the task is completed according to the human’s satisfaction indicated by their utterance this round or if a certain end action is taken. This completion check is domain and task dependent. If the task is not completed, then whole pipeline is repeated for another round of conversation.
IV Experiments
In our experiments, we show that our method enables the robot to perform goal inference to estimate the human’s preference, and then accomplish that task according to it. We conduct experiments for various tasks and human preferences in a text based Grocery Shopping agent domain and an AI2Thor robot simulation domain [27]. We compare our method with two ablation baselines in both of these domains. We also run an isolated inference module experiment on a multiple choice question dataset to see the accuracy of our proposed inference method and to compare the performance between two models of different number of parameters.
IV-A Isolated Inference Intrinsic Evaluation
We check the performance of our proposed goal inference method by testing on a multiple choice dataset [28]. We also compare the performances between using Llama3 8B Instruct and Llama3 70B Instruct models on Hugging Face [29], and check the accuracy with having the “Unspecified Goal” in the goal list. Only open models such as Llama can be used for the inference module because all the logits are available, not just the ones that correspond to the output of text generation.
Each question from the multiple choice dataset has 5 choices. With GPT-4o-mini, we generate four rephrasings for each choice. These 20 choices are the goals in the goal list. For the with “Unspecified Goal” comparison, the “Unspecified Goal” is in the goal list, so there are 21 goals. The sum of the logits that correspond to the words of the original correct choice is taken. The LLM is given the multiple choice question, and each of the 21 goals takes a turn as the “true goal”. If the goal with the highest probability is any of the rephrasings that are derived from the original correct choice, then it is counted as correct. If the most likely goal is any of the rephrasings from the other choices or the “Unspecified Goal”, then it is incorrect.
Figure 3: Isolated Inference Experiments comparing the performance of the Llama3 Instruct 8B vs Llama3 Instruct 70B on only the inference module of the pipeline. Overall, the 8B model has the better performance. The addition of the “Unspecified Goal” does not significantly impact performance.
As seen in Figure 3, we see overall that the 8B Instruct model performs better for both the “Unspecified Goal” not present (94.4%) and “Unspecified Goal” present in the goal list (93.2%) compared to the 70B Instruct model (88.3% and 86.4% respectively). There tiny drop in performance for the “Unspecified Goal” present in the goal list is insignificant, and can be used for determining when to add goals without large impacts on performance. For the remaining experimental results, we use the Meta Llama3 8B Instruct model for inference.
IV-B Experimental Setup
IV-B1 Synthetic Conversation Generation
For our experiments, we simulate human responses with an LLM by conditioning it with a “human profile” at the beginning of the conversation. Both the robot query and the human utterance is generated with GPT4o-mini [30]. All experiments are set at a conversation bound of 20 rounds.
For each round of conversation, the robot generates a question given the robot task description and a transcript of the conversation so far with the human. For the Grocery Shopping agent domain: the task description is “You are a shopping agent that is supposed to make purchases for the human. Your task is to identify a shopping basket that matches the human’s preferences”. For the robot domain: the task description is to “interact with objects in the environment to accomplish the human’s preferences”.
For each round of conversation, given the generated robot query, the human generates a human utterance. We simulate the human with an LLM conditioned on a user human profile consisting of a description of a human’s goal/preference. The human profiles that we do experiments on for the Grocery Shopping domain: five profiles on differing levels of specificity for human profile for gathering ingredients for a chocolate flavored cake (generic chocolate cake, gluten allergy, servings, indecision between flavors, and extra celebratory toppings) along with five other profiles for various meal options are: Italian, SuperBowl, organic, diets for those with anemia. The human profiles for the AI2Thor robot domain experiments are: gathering ingredients for a breakfast sandwich, putting away food, putting away electronic devices, moving valuables into a safe, and gathering cleaning supplies in bathroom.
IV-B2 Inference Module
We can sum together the LLM logits that correspond to tokens of the human utterance when the LLM is given the entire transcript of the conversation and a true goal, $\hat{g}$. Each of the goals in the goal list takes a turn as the “true goal”. Then we convert the log likelihoods to probabilities.
IV-B3 Action Module
In each round of conversation, the action module is responsible for generating a plan and taking actions according to that plan. For our experiments, we use GPT4o-mini as a planner. This can be easily substituted with other planners.
For our experiments, there is a couple of steps to generate a LLM plan. First, there is a LLM generation for a list of applicable objects/search terms based on the most likely goal(s). Then a possible actions list is constructed based on affordances of relevant objects. There is another LLM generation to generate an action plan for given the most likely goal(s) and the possible actions list. If the action plan is not possible, then the LLM needs to regenerate a different plan.
The possible action types for the Grocery Shopping experiment: “search inventory”, “add item to cart”, “remove item from cart”, “buy basket”. The “search inventory” function uses the NLTK package [31] and a Kaggle Grocery Store inventory dataset [32] to implement a simple embedding search by similarity, and retrieves a single most similar item. The shopping basket/cart is represented by a dictionary, so the add item and remove item functions are just dictionary manipulation functions. The “Buy Basket” function is just where the contents of the cart dictionary, total price of the cart are printed out, along with the message “the cart is now purchased”.
The full list of actions that we can use from AI2Thor for the robot domain is Open, Close, Pickup, Put (for each of the the different receptacles that are viable for the object), Push, Pull, Toggle On/Off, Fill/Empty, Slice, Cook, Break, Dirty/Clean. For the generated plan, the robot takes the actions in sequence. If an action in the plan fails, we “undo” the steps that we have taken in the action plan, and then we generate a new action plan. If the whole action sequence is acted out by the robot successfully, then it is added to a successful action transcript list. This successful action transcript list is used to help with the “undo” action and serves as a record of the entire history of successfully taken actions for the task. There is no pre-existing “undo” action in the simulation, we implement it by resetting the environment and then take all the actions in the successful action transcript list.
For taking the actions, the robot uses the “Teleport” function to “Interactable Positions” for specific objects in between each of the actions for the generated action sequence. For the Pickup action, the robot is implemented to also do object retrieval actions such as opening all the current parent receptacles of the object before picking the object up. For the Put action, the robot is implemented to open all parent receptacles of the target receptacle of the object. For kitchen environment, specially linked object pairings such as paired StoveBurner to specific StoveKnob objects information is given to Toggle On/Off actions.
For the Grocery Shopping domain, the task is completed and the conversation ends after the “buy basket” action that is called. For the Robot domain, it is after the “task completed” action that is called. The maximum number of conversation rounds is capped at 20. The current state of the cart or successful action history is given to evaluation.
IV-B4 Goal List Evaluations
We check the performance of the Goal Management module through checking if the proposed goal list is reasonable every round, and if the goals that are removed every round are reasonable as well.
IV-B5 Ablation Baselines
We compare our pipeline with two ablation baselines, the No Goals Baseline and the No Inference Baseline. The No Goals Baseline tests a version of the pipeline that does not use or keep track of goals. The action module generates an action sequence based on the current round human utterance and previous transcript instead of a most likely goal. The No Inference Baseline tests a version of the pipeline that does not use our inference method to calculate the log likelihoods for goals. Instead, additional LLM prompts are used instead to get the most likely goal for action planning and least likely goal for goal removal given the goal list.
IV-B6 LLM Evaluations
We conduct LLM evaluations for our experiments. We evaluate the quality of the generate goal lists. The Goals Reasonable score is assigned out of a overall score of 5, for whether the goals are reasonable given the human utterances, goals information, and high level task for each round of conversation. The Goals Removed Reasonable score is assigned out of a overall score of 3, for whether the goals removed each round are reasonable given the human utterances, goals information, and high level task for each round of conversation.
We also evaluate the outcomes. For the Grocery Shopping Domain, the cart score is assigned out of a score of 3, given the cart, the task of the robot, and the human profile. For the Robot Domain, the successful action transcript for the robot experiment is assigned out of a score of 3, given the final successful action transcript, the task of the robot, and the human profile.
IV-C Results
IV-C1 The Goal List Evaluations
For the Grocery Experiments with our pipeline, the Goals Reasonable score is 97.17% ±$\pm 0.06\%$ and the Goals Removed Reasonable score is 99.09% ±$\pm 0.37\%$. For the Robot Experiments, the Goals Reasonable score is 95.82% ±$\pm 0.48\%$ and the Goals Removed Reasonable score is 93.2% ±$\pm 1.6\%$. For both of these scores, our pipeline does well.
IV-C2 The Outcome Ablation Evaluations
Figure 4: Shopping Domain Ablation Baseline comparisons for LLM evaluations with multiple trials over 10 different human profiles. Our pipeline does better in comparison against a No Goals Baseline that conditions actions on the full dialog. Our pipeline also does better in comparison against a No Inference Baseline that uses a LLM to find the most-likely goal.
Figure 5: Robot Domain Ablation Baseline comparisons for LLM evaluations over multiple trials of five different human profiles. Our pipeline does better in comparison against a No Goals Baseline that conditions actions on the full dialog. Our pipeline also does better in comparison against a No Inference Baseline that uses a LLM to find the most-likely goal.
We compare our method with two ablation baselines for both of our experiment domains. As shown in Figure 4 and Figure 5, our method performs better than both the No Goals Baseline and the No Inference Baseline for both domains. These evaluations are performed by LLM scoring over 5 trials.
V Discussion and Conclusion
We introduced a language-assisted framework for open-ended goal inference that allows for great flexibility in tasks where human preferences are unclear or challenging to specify. We have shown how Language Models can be leveraged to easily edit the support of possible human goals and maintain beliefs over them. We demonstrated that our pipeline can efficiently propose and add new goals based on conversations with the human, and remove goals that are deemed unlikely, undesirable, or unsafe. We also see from our isolated inference evaluation, that overall that the 8B Instruct model performs better compared to the 70B Instruct model.
Future work involves conducting human user studies and see if the LLM evaluations are reflected. While our method shows promise, it currently relies on synthetic conversations generated by LLMs. Future work should validate these results with real human-robot interactions. Other future work can investigate further into less constrained interactions where the human and the agent can take equal roles in providing information, exploration, and inquiry. Our pipeline can also be combined with VLMs, to enable for extracting other representations and amounts of information for both the human and the agent. Our implementation also does not assume that the human needs to answer optimally, and future work can explore cases where the human may be distracted and dividing their attention amongst multiple things, or cases where the human preferences may be slightly altered due to questions asked or persuasion by the agent.
VI Acknowledgements
We thank Jacob Andreas, Phillip Christoffersen, Pinar Ozisik, Belinda Zou Li, Michelle Li, and Andi Peng for helpful conversations. We also thank the reviewers from the HRI-GenAI 2024 Workshop at RSS for their feedback. Rachel Ma was supported by the MIT Presidential Fellowship. Dylan Hadfield-Menell was funded by a Schmidt Sciences AI2050 early career researcher fellowship.